OpenStudy (anonymous):

NEED ALG 2 HELP ASAP PLEASE!

OpenStudy (anonymous):

Here's two of the options as well as the equation.

OpenStudy (anonymous):

@math92130 @mathmale @kropot72

OpenStudy (aum):

Those are the only two choices?

OpenStudy (aum):

2^y = 5 We know 2^2 = 4 and 2^3 = 8. So y has to be somewhere between 2 and 3, but closer to 2. When x = 5, neither of the two graphs has y a little bigger than 2. There must be a couple of more choices in the answers.

OpenStudy (anonymous):

No sorry, let me post the other answers

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@zepdrix

OpenStudy (mathmale):

First, make certain you have all four possible answer choices in front of you. I was a bit thrown off because you initially posted only two of the four possible choices. Let's actually solve the given equation, \[2^y=5\] Please take either the common log (log) or the natural log (ln) of both sides of this equation. What do you get?

OpenStudy (mathmale):

Write out the new equation that these actions give you.

OpenStudy (anonymous):

log(2) 5 = y @mathmale

OpenStudy (anonymous):

@mathmale @zepdrix

OpenStudy (mathmale):

Yes, that's correct. Now, which of the graphs shows a value of 5 for x and a value of ln (to the base 2) of 5 for y?

OpenStudy (mathmale):

Please note that if you are using the "log to the base 2," that's fine, but you could just as well use ln or log; your final answer will be the same.

OpenStudy (mathmale):

Hints: Each of the graphs has a dot on it. The dot represents a point (x,y) on the logarithmic graph. When x=5, what value would you expect y to have? Use a calculator, if need be, ot find that y-value, and then decide which of the four given graphs is the correct one.

OpenStudy (anonymous):

wait okay let me think sry

OpenStudy (anonymous):

When x = 5, y = 2? @mathmale

OpenStudy (mathmale):

Roughly 2. Yes. So, which graph would give you approx. this same info?

OpenStudy (anonymous):

@mathmale the 4th!

OpenStudy (mathmale):

Sorry for my slow response. Yes, that 4th graph is the only one that makes sense. Nice work!

OpenStudy (anonymous):

@mathmale thanks so much! do you mind helping me on one last problem?

OpenStudy (mathmale):

Go ahead and post it (separately from this one, please).

OpenStudy (mathmale):

@chris! I don't see your new question posted yet.

OpenStudy (anonymous):

@mathmale i'll post it and tag you! thanks :)